# Exploring wetting dynamics of water nanodroplet on nano-textured surfaces: atomistic insights

### v1

@@ 3.3 Effect of surface roughness on static contact angle
Following the fact that the introduced roughness on Ti surfaces, Surf-2, 3, 4, and 5 changes the wetting conditions as earlier stated, classic wetting theories such as the Wenzel and Cassie-Baxter theories are used to predict the CA on the surfaces. The results are then compared with the MD numerical simulation to study the effect of the surface roughness on static CA on Ti surfaces. More details on the classical theories can be found in the ”Supporting Information 1.2”.
Fig. 7 shows the comparison of MD simulations in this study with those predicted by classical wetting models, specifically the Wenzel and Cassie-Baxter models, for Ti with nanometric surface reliefs. The flat Ti surface (Surf-1) serves as a baseline for understanding the wetting behavior without any influence of roughness. The CA observed in the MD simulations for this surface is close to the intrinsic Young contact angle. This value sets the stage for comparing how roughness modifies the wettability of the surface.
The MD simulations show an increase in the static CA for Surf-2, 3, and 4 compared to the flat surface (Surf-1). This behavior shows that the droplet conforms to the surface roughness, increasing the actual contact area between the nanodroplet and the nanorough Ti surface. As a result of this, the wetting properties of the surfaces changed from a hydrophilic state to a hydrophobic state. The Wenzel model also predicted a higher CA compared to the flat surface, thus confirming the effect of the introduced relief. However, there is a disagreement between the MD calculations and the Wenzel model, as the measured CA in MD is greater than the CA predicted by the Wenzel model, with an observed deviation of about $15^{\circ}$, $17^{\circ}$ and $21^{\circ}$. Even though Surf-2, 3, and 4 do not entirely agree with the quantitative predictions of the Wenzel model, the qualitative assumptions of the model can still apply when the droplet fills the textured surfaces and no air pockets are present.
@@
As shown earlier for Surf-5, the water nanodroplet does not fully penetrate the surface roughness but instead rests on top of it, trapping air pockets underneath, thereby increasing the CA. Therefore, the CA observed in the MD simulations for Surf-5 is higher than that of the other four surfaces. The Cassie-Baxter model predicts an even higher CA than the MD results, indicating partial agreement between the model and the MD simulations. This partial agreement suggests that Surf-5 exhibits characteristics of the Cassie-Baxter regime, where the reduced solid-liquid contact area results in an increased CA and amplified hydrophobic effect.
@@
Fig. 9 illustrates the relationship between the water-solid interaction energy and the roughness factor (r) of Ti surfaces. This interaction energy, defined as the total van der Waals interaction between water molecules and Ti surfaces is important in understanding the behavior of nanodroplet in contact with varying surface roughness on Ti. For the flat surface (Surf-1), the interaction energy is positive, which indicates a net repulsive force between water molecules and the flat Ti surface leading to weaker adhesion properties with nanodroplet due to smaller surface area for interaction.
In Surf-2, the interaction energy is negative. This sharp drop reveals strong adhesive interactions, which lead to wetting of the rough Ti surface by the nanodroplet. This phenomenon is due to the increased surface area and roughness that increased the Van der Waals interactions between the water molecules and the rough Ti surface. This further gives more validation to why the spreading radius of Surf-2 is significantly higher than the other surfaces as previously discussed. Surf-3 shows a slight reduction in the magnitude of interaction energy. This demonstrates that the highly attractive interactions noted in Surf-2 are slightly reduced as roughness continues to increase.
In Surf-4, the interaction energy becomes less negative, indicating weaker interactions. This trend continues with Surf-5, where the interaction energy is even less negative compared to Surf-4. This demonstrates weaker interactions, leading to non-wetting behavior. These observations show that as the roughness increases, the attractive interactions between the nanodroplet and the Ti surface diminish, making the surface more hydrophobic. This occurs because the nanodroplet does not fully conform to the rough surface, resulting in fewer contact points for Van der Waals interactions. Further increasing the roughness reduces the magnitude of these attractive forces. This finding is consistent with trends in the literature, which show that when the droplet state transitions from the Wenzel state to the Cassie state, the interaction energy decreases sharply [32, 33]. Therefore, understanding and controlling surface roughness can play an important role in tailoring the interfacial properties of Ti surfaces for specific applications. Future research could focus on a deeper investigation of the transition points between different wetting states for a more comprehensive understanding.
@@ 4 Conclusions
MD simulations have been used to investigate the wetting behaviors of a water nanodroplet on flat and structured Ti solid surfaces with different roughness scales. Four different roughness scales have been considered to study both static and dynamic wetting phenomena of water nanodroplets on Ti surface. The static contact angle obtained for the flat Ti surface agrees with the experimental observation. Additionally, a lower spreading radius of water nanodroplets has been observed on Ti surfaces with a high roughness scale compared to other rough surfaces showing the direct correlation between the roughness and the spreading behavior. The spreading process is affected by the ratio of the structure period to the nanodroplet diameter, especially when these values are comparable. The results of the dynamic contact angle of water nanodroplets on the flat Ti surface and those of the rough surfaces showed a gradual decrease over time, illustrating an evolution toward a hydrophobic state. The degree of roughness and the specific dimensions of the surface grooves determine these wetting processes. Surfaces with larger grooves and spacing tend to support more extensive spreading, while those with narrower grooves and higher roughness limit spreading. Our MD simulations further validate this claim even when their results deviate from the data given by the available phenomenological models. It should be noticed that the decrease in the interaction energy facilitates transitions from the Wenzel state to the Cassie state. Overall, we believe that these simulations have improved our understanding of how Ti surface relief can be designed for many applications, in particular for anti-viral surface treatment and other biomedical applications.

### v2

@@ 3.2 Static contact angle of nanodroplet on Ti surfaces
Several measurable macroscopic properties, such as CA, wettability, and surface roughness, are compared with the simulation results to establish consistency. It is, however, important to note that, direct comparison between MD simulations and experimental results faces certain challenges, particularly due to the differences in scale. The simulations in this study deal with nanodroplets, whereas most experimental studies use larger microscale droplets.
We note that the behavior of nanodroplets, like the one of nanoparticles, is more affected by their surface tension and can reveal remarkable features that are still far from being completely understood. Although the qualitative trends such as the transition from Wenzel to Cassie-Baxter states are similar across scales, quantitative discrepancies may arise due to size effects. Furthermore, experimental surfaces often have imperfections and irregularities that are not captured in idealized simulation models, potentially causing deviations in observed CA or droplet dynamics. Additionally, environmental factors such as temperature and pressure, which are tightly controlled in simulations, may vary in experimental conditions, introducing further variability.
@@ 3.3 Effect of surface roughness on static contact angle
Following the fact that the introduced roughness on Ti surfaces, Surf-2, 3, 4, and 5 changes the wetting conditions as earlier stated, classic wetting theories such as the Wenzel and Cassie-Baxter theories are used to predict the CA on the surfaces. The results are then compared with the MD numerical simulation to study the effect of the surface roughness on static CA on Ti surfaces. More details on the classical theories can be found in the "Supporting Information 1.2".
Fig. 7 shows the comparison of MD simulations in this study with those predicted by classical wetting models, specifically the Wenzel and Cassie-Baxter models, for Ti with nanometric surface reliefs. The flat Ti surface (Surf-1) serves as a baseline for understanding the wetting behavior without any influence of roughness. The CA observed in the MD simulations for this surface is close to the intrinsic Young CA. This value sets the stage for comparing how roughness modifies the wettability of the surface.
The MD simulations show an increase in the static CA for Surf-2, 3, and 4 compared to the flat surface (Surf-1). This behavior shows that the droplet conforms to the surface roughness, increasing the actual contact area between the nanodroplet and the nanorough Ti surface. As a result of this, the wetting properties of the surfaces changed from a hydrophilic state to a hydrophobic state. The Wenzel model also predicted a higher CA compared to the flat surface, thus confirming the effect of the introduced relief. However, there is a disagreement between the MD calculations and the Wenzel model, as the measured CA in MD is greater than the CA predicted by the Wenzel model, with an observed deviation of about $15^{\circ}$, $17^{\circ}$ and $21^{\circ}$ respectively. Even though Surf-2, 3, and 4 do not entirely agree with the quantitative predictions of the Wenzel model, the qualitative assumptions of the model can still apply when the droplet fills the textured surfaces and no air pockets are present.
As shown earlier for Surf-5, the water nanodroplet does not fully penetrate the surface roughness but instead rests on top of it, trapping air pockets underneath, thereby increasing the CA. Therefore, the CA observed in the MD simulations for Surf-5 is higher than that of the other four surfaces. The Cassie-Baxter model predicts an even higher CA than the MD results, indicating partial disagreement between the model and the MD simulations. Despite this disagreement, Surf-5 exhibits characteristics of the Cassie-Baxter regime, where the reduced solid-liquid contact area results in an increased CA and amplified hydrophobic effect.
It is worthy of note that the observed deviations between the MD simulation CAs and the predictions of classical wetting models can be due to several factors that highlight the difficulties of wetting at the nanoscale. The classical models typically consider macroscopic droplets, while this study focuses on nanodroplets, where surface effects and molecular interactions become more pronounced. The increased surface-to-volume ratio of nanodroplets can lead to different wetting behavior compared to larger droplets, as the relative contributions of surface energy and line tension become more significant. Furthermore, the dynamic nature of wetting, with evolving droplet shapes and molecular rearrangements, may not be fully captured by static models like Wenzel and Cassie-Baxter. These models provide a snapshot of the wetting state at a particular time but do not account for the time-dependent evolution of the droplet shape and CA.
@@
Fig. 9 illustrates the relationship between the water-solid interaction energy and the roughness factor (r) of Ti surfaces. This interaction energy, defined as the total Van der Waals interaction between water molecules and Ti surfaces is important in understanding the behavior of nanodroplet in contact with varying surface roughness on Ti. Specifically, we evaluated the water-solid interaction energy for each surface to quantify how roughness influences wetting behavior. In addition to several analytical models that assumes that only shape and the slope of the pillar edge are responsible for the unusual water droplet spreading, we showed that surface energy effects are in line with previous experimental and theoretical observations [65, 66, 67]. In particular, these calculations help to clarify the correlation between surface roughness and the observed transitions between the Wenzel and Cassie-Baxter wetting regimes. This approach will provide a more comprehensive understanding of how nanostructured surfaces affect energy dissipation during the wetting process.
For the flat surface (Surf-1), the interaction energy is positive, which indicates a net repulsive force between water molecules and the flat Ti surface leading to weaker adhesion properties with nanodroplet due to smaller surface area for interaction. In Surf-2, the interaction energy is negative. This sharp drop reveals strong adhesive interactions, which lead to wetting of the rough Ti surface by the nanodroplet. This phenomenon is due to the increased surface area and roughness that increased the Van der Waals interactions between the water molecules and the rough Ti surface. This further gives more validation to why the spreading radius of Surf-2 is significantly higher than the other surfaces as previously discussed. Surf-3 shows a slight reduction in the magnitude of interaction energy. This demonstrates that the highly attractive interactions noted in Surf-2 are slightly reduced as roughness continues to increase.
In Surf-4, the interaction energy becomes less negative, indicating weaker interactions. This trend continues with Surf-5, where the interaction energy is even less negative compared to Surf-4. This demonstrates weaker interactions, leading to non-wetting behavior. These observations show that as the roughness increases, the attractive interactions between the nanodroplet and the Ti surface diminish, making the surface more hydrophobic. This occurs because the nanodroplet does not fully conform to the rough surface, resulting in fewer contact points for Van der Waals interactions. Further increasing the roughness reduces the magnitude of these attractive forces. This finding is consistent with trends in the literature, which show that when the droplet state transitions from the Wenzel state to the Cassie state, the interaction energy decreases sharply [44, 45, 67]. Therefore, understanding and controlling surface roughness can play an important role in tailoring the interfacial properties of Ti surfaces for specific applications.
@@ 4 Conclusions
This study provides atomic-level insights into the wetting mechanisms of water nanodroplets on both flat and nano-textured titanium surfaces, using molecular dynamics simulations, with implications for medical implants. Our results demonstrate a clear correlation between surface roughness and wettability, with higher roughness promoting a transition from Wenzel to Cassie-Baxter wetting regimes. The simulation results also show that the introduction of surface roughness significantly influences the spreading behavior of the nanodroplet, with larger grooves promoting extensive spreading and narrower grooves limiting it. Furthermore, we observed the damping of droplet vibrations and a transition to a hydrophobic state on both flat and rough surfaces, driven by atomic bonding mechanisms specific to Ti surfaces. The MD simulations reveal interaction energy trends that show strong adhesive interactions on rough surfaces, leading to increased hydrophobicity as roughness further increases. The discrepancies between simulation contact angle and classical wetting models shows the importance of considering nanoscale effects and molecular-level interactions in understanding and predicting wetting phenomena. These findings have significant implications for the design and optimization of titanium-based biomaterials, as well as for the development of advanced materials with tailored wetting properties for various applications, including self-cleaning surfaces and microfluidic devices.
Supporting Information
The Supporting Information is available free of charge at [Link].
Calculated structural properties of titanium, snapshots of water nanodroplets on flat titanium surface, and classical wetting models, including the Wenzel and Cassie-Baxter models [PDF].